f:X to Y f(x)=sinx+cox+2root2 is invertible then X to Y is

f:X to Y
f(x)=sinx+cox+2root2 is invertible then X to Y is 

Grade:9

1 Answers

Arun Kumar IIT Delhi
10 years ago
Hello Student,
\\f(x)=\sqrt2 *sin(x+{ \pi \over 4})+2\sqrt2 \\$or $f(x)=\sqrt2 *cos(x+{ \pi \over 4})+2\sqrt2 \\=>Y=[\sqrt2,3\sqrt2] \\=>X=[{-3\pi \over 4},{ \pi \over 4}]$ or $[{\pi \over 4},{ 5\pi \over 4}]

Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty